OpenStudy (anonymous):

In a function, how will the y-values be different if x is negative or positive? A. The y-values will decrease if x is negative and increase if x is positive. B. The y-values will increase if x is negative and decrease if x is positive. C. There will be no difference. If x increases, y-values will increase. D. You can't tell without knowing the function rule.

OpenStudy (anonymous):

I think it is D. am I right

OpenStudy (lukebluefive):

Consider two functions: \[y = x + 1\] \[y = -x - 1\] Ask these questions for each function and see which one holds with both.

OpenStudy (lukebluefive):

Put more simply, if you increase x in both functions, will y respond the same way in both functions?

OpenStudy (anonymous):

yes

OpenStudy (lukebluefive):

For the first function, let x = -2: \[y = -2 + 1 = -1\] Then let x = 2: \[y = 2 + 1 = 3\] So, if x is positive, y is usually positive, and if x is negative, y is negative. Increasing x leads to an increase in y. How about for the second function? Let x = -2: \[y = -(-2) - 1 = 2 - 1 = 1\] And then let x = 2: \[y = -2 - 1 = -3\] Here, it is the opposite! If x is positive, y is negative, and if x is negative, y is positive. Furthermore, increasing x decreases y. Therefore, what can we say about these two functions? Which of the answers holds?

OpenStudy (anonymous):

a

OpenStudy (lukebluefive):

For the first function A & C apply. For the second function only B applies. What's the only answer that's consistent with this result?

OpenStudy (anonymous):

it would be D